OpenStudy (waterbird):

cosϴ=-√3 sin(-ϴ) HELP NOW PLEASE!!!

OpenStudy (mww):

if you wish to solve for theta best thing to do is to generate the tan ratio \[\frac{ \sin \theta }{ \cos \theta } = \tan \theta\] Also note that sin is an odd function so \[\sin(- \theta) = - \sin(\theta)\] Use the combination to generate an equivalent equation in tan.